OpenStudy (anonymous):

A tree is 3,212 years old. How can you express this age in scientific notation with the LOWEST level of precision? a. 3x10^3 b. 3.2x10^3 c. 3.21x10^3 d. 3.212x10^3 I think b is the answer, but if I'm wrong, I'd like someone to correct me.

OpenStudy (anonymous):

i believe that the answer is a

OpenStudy (anonymous):

Would you know how or why? Im just trying to have a better understanding.

OpenStudy (anonymous):

well, 3000 is less precise than 3200 :)

OpenStudy (anonymous):

3*10^3 = 3*1000 = 3000 3.2*10^3 = 3.2*1000 = 3200; 3200 is closer to 3212, thus more precise

OpenStudy (anonymous):

I never thought of that, I thought it would have to stay in the 200 area. But it makes a lot more sense now. Thank you!

OpenStudy (anonymous):

you're welcome :)

OpenStudy (anonymous):

help